# Aging properties of the voter model with long-range interactions

### v1

@@ I Introduction
On the r.h.s., the former is a time-translation invariant contribution provided by fast degrees of freedom that locally equilibrate with the thermal bath. The latter term represents what is left over, namely the genuinely off-equilibrium effects. A physical intuition can be more easily achieved for scalar systems, i.e. $S_{i}=\pm 1$, to which we will always restrict in the following. In this case the post-quench evolution is characterized by formation and growth of ordered domains of characteristic size $L(t)$, whose interior is representative of the two, symmetry-related, pure states at low-temperature, a phenomenon dubbed coarsening. Then, fast degrees involved in the first term on the r.h.s. of Eq. (2) are the so-called thermal islands, namely the ephemeral reversal of spins in small regions, of the size of the equilibrium coherence length, well inside the bulk of otherwise ordered domains. On the other hand $G_{ag}$ encodes the non-equilibrium motion of the domain walls.
$G_{eq}$, being an equilibrium quantity, is generally quite well understood. In the following, therefore, we will be only interested in $G_{ag}$, which can be conveniently isolated, for instance, by considering zero-temperature quenches, because $G_{eq}\equiv 0$ in this case. This is the situation we will consider in this paper and, for this reason, we will set $G_{ag}=G$ in the following, which lightens the notation. This quantity has been extensively studied [2, 3, 4, 1, 5, 6, 7, 8, 9, 10, 11] in systems with short-range interactions. Here it is generally found that $G$ obeys a scaling form
@@
where $g$ is a scaling function. This implies
@@
In this respect, the possibility to study solvable models is mostly desired. A possibility is offered by the voter model, whose dynamical equations can be closed in any dimension [32, 33] and for every kind of interaction [34, 35]. Of course, despite being ferromagnetic in spirit, the model is not fully representative of physical magnets, because its kinetic rules generally lack detailed balance. However it reduces to the Ising model in space dimension $d=1$ with NN and is interested by a coarsening phenomenon also with extended interactions and in $d=2$ as well. For $d\geq 3$, instead, coarsening is interrupted.
@@
The voter model [36, 37, 38, 39, 40, 32, 33, 41, 42, 43, 44, 45] was developed and popularized in the framework of ecological systems [36] as an invasion model, and later on in opinion dynamics [37], although some preliminaries were already investigated to study the spread of genetic correlations [46, 47, 48]. It can be described in this way: at each vertex of a lattice, an agent can adopt one of two options, typically represented as “left” and “right”, which can equivalently be encoded in a binary variable $S_{i}=\pm 1$, also called spin for the manifest similarity with the Ising model [49, 50, 51, 52, 53, 54, 55]. In its original definition an agent’s choice is influenced by a randomly selected neighboring agent, therefore with the probability of spin-flip being proportional to the fraction of opposite-spin neighbors.
@@
Over the years, numerous variants of the original model have been proposed to adapt it to explain various situations [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 53, 67]. In such a context, a long-range variant of the voter model has been studied, where an agent at a lattice vertex adopts the opinion of one located at a distance $r$ from him, selected randomly with a probability $P(r)\propto r^{-\alpha}$ [34, 35]. Various regimes of $\alpha$ were analyzed, revealing a rich and diverse structure. For $d<3$ a critical value $\alpha_{c}(d)$ of $\alpha$ exists, such that for $\alpha>\alpha_{c}(d)$ the model behaves similarly to the NN case, with $L(t)$ growing as $\sqrt{t}$ until consensus, namely perfect ordering, is reached. It has been found that $\alpha_{c}(1)=3$ and $\alpha_{c}(2)=4$. Additionally, the correlation function exhibits dynamical scaling behavior, albeit with logarithmic corrections at large distances in the case $d=2$. Such corrections are due to the fact that $d=2$ is a critical dimensionality for the model. As $\alpha$ decreases below a certain threshold, there is the formation of partially ordered stationary states, which in the original NN model are only present for $d>2$. For $\alpha$ small enough, the mean-field behavior ($\alpha=0$) is resembled.
@@
In this paper we study this model in $d=1$, focusing on the aging properties that were not addressed before. Using both analytical and numerical methods to solve the exact closed dynamical equations (see, for $G$, Eq. (9)), we calculate the two-time correlations and investigate their properties across various intervals of the parameter $\alpha$. A very brief and non-exhaustive summary of our results follows: For $\alpha>3$, the behavior is akin to one of the NN model, the autocorrelation function decays with $\lambda=1$ and $G$ obeys the scaling of Eq. (3), where $L(t)$ is a coherence length that can be obtained from the equal-time correlation function $G(r;t,t)$. In the range $2<\alpha\leq 3$ and $1<\alpha\leq 2$, $A(z)$ still displays a power-law decay characterized by $\lambda=1$ and $\lambda=(\alpha-1)^{-1}$, respectively. Scaling of $G$ is still in the form (3), with the additional feature that, only for $1<\alpha\leq 2$, the scaling length is not the coherence length. When $\alpha\leq 1$, the autocorrelation function decays exponentially as in mean-field.
@@ II The model
The factor $2$ appears because there are two ways (on the right or on the left, say) to choose a spin at distance $\ell$ from a reference one. The sum extends up to $N/2$ because we consider a system of $N$ agents with periodic boundary conditions, where $N/2$ is the largest possible distance. From now on we will drop the summation range when sums over $\ell$ will be considered, and it will be understood that $\sum_{\ell}=\sum_{\ell=1}^{N/2}$.
@@ V.1 Coarsening stage
We start by studying the aging properties in the coarsening stage, where we remind that $L\propto t$ in the range of $\alpha$ considered here [34].
@@ V.2 Stationary states
We start noticing that, as already observed, Eqs. (11) and (9) are equal, apart from a trivial factor $2$ and the behavior in $r=0$. This mean that the form
@@ VIII Discussion and conclusions
This paper is the first where the aging properties of the long-ranged voter model have been considered. Even if we tried to provide a rather comprehensive panorama of the behavior of the two-time correlations in the whole range of $\alpha$ values, a lot of questions remain unanswered. Firstly, since a closed equation as Eq. (9) can be written in any spatial dimension, the same topic considered here could be in principle studied in $d>1$. We know [35], that in $d=2$ the one-time properties of the model are quite different from the $1d$ case, so we expect new scenarios to emerge. Also, the response function is another interesting two-time quantity, whose properties are important also in connection with the two-time correlations because of the so-called fluctuation-dissipation relation, whereby a possible definition of an effective temperature springs out. This and other issues remain for future investigations.

### v2

@@ I Introduction
On the r.h.s., the former is a time-translation invariant contribution provided by fast degrees of freedom that locally equilibrate with the thermal bath. The latter term represents what is left over, namely the genuinely off-equilibrium effects. A physical intuition can be more easily achieved for scalar systems, i.e. $S_{i}=\pm 1$, to which we will always restrict in the following. In this case the post-quench evolution is characterized by the formation and growth of ordered domains of characteristic size $L(t)$, whose interior is representative of the two, symmetry-related, pure states at low-temperature (say, with positive and negative magnetization), a phenomenon dubbed coarsening. Then, fast degrees involved in the first term on the r.h.s. of Eq. (2) are the so-called thermal islands, namely the ephemeral reversal of spins in small regions, of the size of the equilibrium coherence length, caused by thermal fluctuations well inside the bulk of otherwise ordered domains. On the other hand $G_{ag}$ encodes the non-equilibrium motion of the domain walls.
$G_{eq}$, being an equilibrium quantity, is generally quite well understood. In the following, therefore, we will be only interested in $G_{ag}$, which can be conveniently isolated, for instance, by considering zero-temperature quenches, because $G_{eq}\equiv 0$ in this case, since thermal island cannot be activated with vanishing temperature. This is the situation we will consider in this paper and, for this reason, we will set $G_{ag}=G$ in the following, which lightens the notation. This quantity has been extensively studied [2, 3, 4, 1, 5, 6, 7, 8, 9, 10] in systems with short-range interactions. Here it is generally found that $G$ obeys a scaling form
@@
where $g$ is a scaling function. Similarly to what is known in equilibrium critical phenomena, this structure reflects the physical fact that, at any time $t$, there is a unique physically relevant length $L$ and that the sole effect of the elapsing of time is the increase of such length. Then, the system is statistically similar to itself at any time, provided that $L$ is used as a unit to measure distances, as it can be seen by the first entry of $g$ in the above equation. The second entry expresses the fact that the passage of time, from $s$ to $t$, can be fully expressed by the fractional increase $L(t)/L(s)$ of the relevant length $L$. Eq. (3) implies
@@
In this respect, the possibility to study solvable models is mostly desired. A possibility is offered by the voter model, whose dynamical equations can be closed in any dimension [31, 32] and for every kind of interaction [33, 34]. Of course, despite being ferromagnetic in spirit, the model is not fully representative of physical magnets, because its kinetic rules generally lack detailed balance. However it reduces to the Ising model in space dimension $d=1$ with NN and is interested by a coarsening phenomenon also with extended interactions and in $d=2$ as well. For $d\geq 3$, instead, there is no asymptotic coarsening.
@@
The voter model [35, 36, 37, 38, 39, 31, 32, 40, 41, 42, 43, 44] was developed and popularized in the framework of ecological systems [35] as an invasion model, and later on in opinion dynamics [36], although some preliminaries were already investigated to study the spread of genetic correlations [45, 46]. It can be described in this way: at each vertex of a lattice, an agent can adopt one of two options, typically represented as “left” and “right”, which can equivalently be encoded in a binary variable $S_{i}=\pm 1$, also called spin for the manifest similarity with the Ising model [47, 48, 49, 50, 51, 43, 52]. In its original definition an agent’s choice is influenced by a randomly selected neighboring agent, therefore with the probability of spin-flip being proportional to the fraction of opposite-spin neighbors.
@@
Over the years, numerous modifications of the original model have been proposed to adapt it to explain various situations [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 51, 64]. In such a context, a long-range variant of the voter model has been studied, where an agent at a lattice vertex adopts the opinion of one located at a distance $r$ from him, selected randomly with a probability $P(r)\propto r^{-\alpha}$  [33, 34]. In these papers the behavior of one-time quantities, specifically the equal time correlation function $G(r;t,t)$ was considered, from which information on the coherence of the system and the approach to the ordered state can be extracted. Various regimes of $\alpha$ were analyzed, revealing a rich and diverse structure. For $d<3$ a critical value $\alpha_{c}(d)$ of $\alpha$ exists, such that for $\alpha>\alpha_{c}(d)$ the model behaves similarly to the NN case, with $L(t)$ growing as $\sqrt{t}$ until consensus, namely perfect ordering, is reached. It has been found that $\alpha_{c}(1)=3$ and $\alpha_{c}(2)=4$. Additionally, the correlation function exhibits dynamical scaling behavior, albeit with logarithmic corrections at large distances in the case $d=2$. Such corrections are due to the fact that $d=2$ is a critical dimensionality for the model. As $\alpha$ decreases below a certain threshold, there is the formation of partially ordered stationary states, which in the original NN model are only present for $d>2$. For $\alpha$ small enough, a behavior akin to mean-field ($\alpha=0$) is recovered.
In this paper we study this model in $d=1$, focusing on the aging properties that were not addressed before. Although, clearly, the results derived here for the two-time quantities discussed above are completely new and can by no means be inferred from the equal time quantities considered in  [33, 34], the present article must be regarded as a companion paper and many useful details can be found in the above references. Anyway, we will also provide the reader with the necessary context to ease the reading.
@@
In this paper, using both analytical and numerical methods to solve the exact closed dynamical equations (see, for $G$, Eq. (9)), we calculate the two-time correlations and investigate their properties across various intervals of the parameter $\alpha$. A very brief and non-exhaustive summary of our results follows: For $\alpha>3$, the behavior is akin to one of the NN model, the autocorrelation function decays with $\lambda=1$ and $G$ obeys the scaling of Eq. (3), where $L(t)$ is a coherence length that can be obtained from the equal-time correlation function $G(r;t,t)$. In the range $2<\alpha\leq 3$ and $1<\alpha\leq 2$, $A(z)$ still displays a power-law decay characterized by $\lambda=1$ and $\lambda=(\alpha-1)^{-1}$, respectively. Scaling of $G$ is still in the form (3), with the additional feature that, only for $1<\alpha\leq 2$, the scaling length is not the coherence length. When $\alpha\leq 1$, the autocorrelation function decays exponentially as in mean-field.
@@ II The model
The factor $2$ appears because there are two ways (on the right or on the left, say) to choose a spin at distance $\ell$ from a reference one. The sum extends up to $N/2$ because we consider a system of $N$ agents with periodic boundary conditions, where $N/2$ is the largest possible distance. From now on we will drop the summation range when sums over $\ell$ will be considered, and it will be understood that $\sum_{\ell}=\sum_{\ell=1}^{N/2}$. With this form of $P(\ell)$ it is clear that the moment of order $n$, $\langle\ell^{n}\rangle\equiv\sum_{\ell}\ell^{n}P(\ell)$, diverges with $N$ for $\alpha\leq n+1$. It can be expected that the moments with $n\leq 2$, representing the normalization of probability, and a characteristic length and length fluctuation related to $P(\ell)$, are relevant to the physical behavior. Therefore, their divergence in the thermodynamic limit, occurring at $\alpha=1,2,3$, respectively, will be accompanied by a marked change of behavior of the model. This is confirmed by the careful analysis of Ref.  [33] and by the results presented below. Such feature, therefore, is a direct consequence of the power-law behavior of $P(\ell)$.
@@
Before proceeding further, let us briefly discuss the idea behind selecting a power-law behavior for the probability $P(\ell)$. Firstly, our focus primarily lies on the system’s behavior for large $t$. Because the domains’ typical size increases over time, only the asymptotic behavior of $P(\ell)$ for large $\ell$ is relevant. Therefore, even more intricate forms of $P(\ell)$, sharing the same asymptotic form $\ell^{-\alpha}$, yield equivalent outcomes. Forms of $P(\ell)$ that asymptotically decay faster than $\ell^{-\alpha}$ (e.g., exponentially) exhibit behavior akin to the case where $\alpha>3$ in the present work, resembling that of the NN model. Conversely, for $P(\ell)$ decreasing slower than $\ell^{-\alpha}$ (e.g., logarithmically), the system likely behaves as in the $\alpha<1$ scenario of the present work, i.e. akin to the mean-field case.
@@ V.1 Coarsening stage
We start by studying the aging properties in the coarsening stage. Let us remind that in Ref. [33] Eq.(11) was explicitly solved, and the computation led to $L\propto t$.
@@ V.2 Stationary states
We start noticing that, as already observed, Eqs. (11) and (9) are equal, apart from a trivial factor $2$ and the behavior in $r=0$. This means that the form
@@ VIII Discussion and conclusions
This paper is the first where the aging properties of the long-ranged voter model have been considered. Even if we tried to provide a rather comprehensive panorama of the behavior of the two-time correlations in the whole range of $\alpha$ values, a lot of questions remain unanswered. Firstly, since a closed equation as Eq. (9) can be written in any spatial dimension, the same topic considered here could be in principle studied in $d>1$. We know [34], that in $d=2$ the one-time properties of the model are quite different from the $1d$ case, so we expect new scenarios to emerge. Also, the response function is another interesting two-time quantity, whose properties are important also in connection with the two-time correlations because of the so-called fluctuation-dissipation relation, whereby a possible definition of an effective temperature springs out. This and other issues, like, e.g., the study of the persistence properties [40], remain for future investigations.